# Plastic Arbor: a modern simulation framework for synaptic plasticity – from single synapses to networks of morphological neurons

### v2

@@ 4 Discussion
In the following, we will briefly discuss the pros and cons of the most prominent neural network simulators, and compare their target use cases with Arbor. With its first version released in the 1980s and widespread usage, NEURON has significantly advanced the understanding of the brain by facilitating the computational study of complex neuronal processes [71]. Nevertheless, as mentioned above, its engine under the hood is outdated when it comes to high-performance computing, especially for network simulations involving detailed neuron models. While CoreNEURON [2, 5] constitutes an approach to address this, it has restricted flexibility and usability due to its dependence on the NEURON environment, and it lacks support for certain important features of NMODL [78]. Further, GENESIS (GEneral NEural SImulation System) [79] is a simulator that has also been used widely for several decades, offering a platform for simulating both small-scale detailed models and large-scale network simulations. However, the use of GENESIS with respect to modern hardware backends is very limited. Regarding alternatives that are more focused on point-neuron simulations, NEST (NEural Simulation Tool) is a widely used simulator designed for large-scale simulations of spiking neural networks. It is particularly known for its scalability and efficiency in simulating large-scale networks. However, its lack of support for multi-compartment neurons has been one of the initial reasons to develop Arbor, which was originally named ‘NestMC’. Brian 2 is another widely used network simulator that enables the definition of models directly via differential equations [4]. Furthermore, Brian 2 enables high flexibility by generating an intermediate abstract code, processed by so-called device interfaces. This allows to seamlessly exchange the underlying numerical backend architecture. While Brian 2 does not yet come with comprehensive support for multi-compartment neurons, a new extension called Dendrify specifically focuses on detailed dendritic morphology and may provide a valuable tool for the investigation of dendritic processing [80]. CARLsim [81] is another framework for the parallelized simulation of large-scale spiking neural networks, but it has been optimized for real-time simulations and neuromorphic hardware implementation. Accordingly, the single-compartment Izhikevich neuron is the most biologically realistic neuron model that is supported by CARLsim. Finally, the recently released simulator EDEN [82] appears to offer high flexibility by supporting NeuroML model descriptions, however, it still lacks support for GPU backends. As opposed to that, the equally new simulator NeuronGPU [83] is optimized for neural network simulations on GPU but does currently not support multi-compartment models.
@@
In summary, each of the existing simulators comes with its own strengths and weaknesses (also see [84, 85, 86]). The decision to use one specific simulator should depend on the particular needs of a research project, the available compute resources, as well as the expertise of the involved researchers. For example, Brian 2 can be considered quite user-friendly for quickly setting up networks of point neurons, NEST exhibits unique performance for very large network simulations, and for NEURON there are many existing implementations of morphological neuron models. Arbor, finally, was designed to easily define networks of morphological neurons, and then to map the internal modeling primitives to available compute resources in an optimized manner. This focus makes Arbor an attractive option, particularly, for researchers who rely on high-performance computing for their simulations with networks of morphologically detailed neurons.
@@
We have cross-validated all of our presented Arbor implementations either with Brian 2 [4] or with one of multiple stand-alone simulators that were custom-developed for previous studies [20, 58, 45]. At the single-synapse level, we did not find any significant differences between the results of the simulators (Figs. 2–5; Supplementary Figs. S1–S3). However, certain differences in the neuron model and in the numerical methods, as well as complex network effects, give rise to certain quantitative deviations in memory recall performance at the circuit level, while the qualitative behavior remained the same (Supplementary Fig. S10; also cf. subsection 3.5 and Supplementary Fig. S4).
@@
Regarding the use of compute resources for a large network with plastic connections, we found that Arbor performs well both in terms of runtime and memory use. Compared to optimized point-neuron simulators, Arbor only uses slightly more resources when computing on CPU, and can even outperform those simulators by computing on GPU (Fig. 8). However, Arbor particularly shines when it comes to simulating networks of multi-compartment neurons, which necessitates almost no additional cost compared to single-compartment neurons (shown by the runtime per spike, Supplementary Fig. S12). This is not entirely unexpected, since Arbor has been designed particularly for networks of multi-compartment neurons.
@@
The benchmarking results notwithstanding, there may be several ways to further improve the performance of Arbor. First, to compute the dynamics of given models, Arbor uses implicit solving methods, which have the advantage of being stable but come at the cost of runtime performance. In specific cases, these methods may be replaced by faster explicit algorithms. Kobayashi et al. [87], for instance, have shown how an explicit method with adaptive time steps and second-order accuracy can serve to avoid heavy memory access, which can be helpful particularly when using GPUs. Second, the introduction of exact point neurons may eliminate the need to simulate a spatial neuron model if this is not needed. Note that meanwhile, a LIF neuron feature has been added to Arbor, although still being in a test stage.
@@
By the final model simulation results that we have presented in Fig. 6, we could gain first new insights into the functional interplay between, on the one hand, network parameters such as the size of a stimulated pattern, and on the other hand, PRP diffusion within neurons. Specifically, the results suggest that the functionality of pattern recall is not very sensitive to the diffusion speed. Nevertheless, if the diffusion is too slow, as expected, all memory functionality will vanish, which also occurs if the pattern size is too small. Moreover, if we increase the size of the dendritic branches, the memory recall capability is impaired (Fig. 6f–h), indicating that additional mechanisms may be needed to obtain a stable memory system. Somehow in contrast to that, a larger cell and dendrite diameter improves memory recall (Fig. 6j–l). By targeting the impact that PRP diffusion within neurons has on the dynamics of large networks, our findings may complement the picture that other studies have drawn of the functional role of spatial PRP dynamics within neurons [88, 89, 90, 91, 92].
@@
In the future, most importantly, the extended Arbor simulation framework can enable researchers to conduct studies that examine the interplay between neuron-internal and network dynamics. Moreover, Arbor allows the implementation of models that include changes in the connectivity structure of networks. Using this together with the tools that we have presented here will enable researchers to also study the interactions between structural and functional plasticity processes at the network level (cf. [93]). It should further be noted that by considering synaptic plasticity processes in neural networks, simulation software contributes essentially to the understanding of biological mechanisms as well as to the development of artificial intelligence applications. In this light, further applications of our extended version of Arbor may target, for example, paradigms of somato-dendritic mismatch error reduction [94], reservoir computing with heterogeneous networks [95, 96], or prototyping and benchmarking of neuromorphic algorithms (cf. [85, 97]).
@@
Due to its modern computing architecture and inherent support of multi-compartment neurons, the Arbor simulator constitutes an important tool for the computational modeling of neuronal networks. Through the newly introduced extensions, enabling to simulate synaptic plasticity, we have increased the range of Arbor’s use cases substantially. Furthermore, we have provided a first demonstration of how the extended framework enables to gain new insights into the functional impact of morphological neuronal structure at the network level. In the future, the Plastic Arbor framework may power a great variety of studies considering synaptic mechanisms and their interactions with neuronal dynamics and morphologies, from single synapses to large networks.
Data availability
@@
The Arbor framework can be installed as described on https://arbor-sim.org (the code can be retrieved, e.g., from [7]). Our simulation code for the different subsections is referenced in Table 7 below. Note that since Arbor is still under development, some parts of our model implementations may be subject to changes when using later Arbor versions, however, we intend to keep the model implementations updated along with the Arbor core code.
Due to the ongoing development of the Arbor core code, we ran our extensive simulations with different Arbor versions. For the simulation results presented in subsections 3.1–3.3 as well as 3.5–3.6, we used Arbor version 0.9.1-dev-2f4c325 (for instructions on how to install this exact version, see the README.md file that comes with our model code [67]). For the simulation results of subsection 3.4, we used the development branch of pull request #2226 with the state of commit f0e456d due to a necessary fix in the diffusion processing; the branch shall be merged upon publication of this article (similarly, Arbor still considers diffusion constants in non-standard units, scaled by a factor of , which will be corrected as well). For the benchmarking results in subsection 3.7 we used Arbor version 0.10.0. Note that while the mentioned Arbor versions are by now slightly outdated, we ensured that our extensive software tests from [67, 98] passed with the used versions including the release version 0.10.0, as well as with the currently latest version 0.10.1-dev-b8b768d, which strongly indicates that the simulation results will remain the same.
@@ Acknowledgments
We would like to thank the other Arbor developers for their support and for their dedication to continuously improve the software. We would further like to thank Silvio Rizzoli and Arash Golmohammadi for helpful discussions.
Funding
This work was supported by the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) through grants SFB1286 (C01, Z01) and TE 1172/7-1, as well as by the European Commission H2020 grants no. 899265 (ADOPD) and 945539 (HBP SGA3).
@@
Conceptualization: JL, SSc, CT; Data curation: JL, SSc, SSh; Formal Analysis: JL, SSc, SSh; Funding acquisition: SSc, CT; Investigation: JL, SSc, SSh, CT; Methodology: JL, SSc, SSh, CT; Project administration: CT; Resources: TH, CT; Software (Arbor core): JL, SSc, TH, FB (with larger contributions by TH and FB); Software (sections 3.1–3.2): JL, SSc; Software (section 3.3): SSc, FB; Software (section 3.4): SSh; Software (sections 3.5–3.7): JL; Supervision: JL, SSc, CT; Validation: JL, SSc, SSh, TH, FB; Visualization: JL, SSc, SSh, TH; Writing – original draft: JL, SSc, SSh, TH, FB; Writing – review & editing: JL, SSc, CT.

### v3

@@ 4.1 Discussion and outlook
In the following, we will briefly discuss the pros and cons of the most prominent neural network simulators, and compare their target use cases with Arbor. With its first version released in the 1980s and widespread usage, NEURON has significantly advanced the understanding of the brain by facilitating the computational study of complex neuronal processes [72]. Nevertheless, as mentioned above, its engine under the hood is outdated when it comes to high-performance computing, especially, for network simulations involving detailed neuron models. While CoreNEURON [2, 5] constitutes an approach to address this, it has restricted flexibility and usability due to its dependence on the NEURON environment, and it lacks support for certain important features of NMODL [82]. Further, GENESIS (GEneral NEural SImulation System) [83] is a simulator that has also been used widely for several decades, offering a platform for multi-scale simulations with particular focus on detailed electrical and chemical interactions. While GENESIS is very limited with respect to modern hardware backends, the MOOSE simulator is developed as its modern successor [84]. Regarding alternatives that are more focused on point-neuron simulations, NEST (NEural Simulation Tool) is a widely used simulator designed for large-scale simulations of spiking neural networks. It is particularly known for its scalability and efficiency in simulating large-scale networks. However, its lack of support for multi-compartment neurons has been one of the initial reasons to develop Arbor, which was therefore originally named ‘NestMC’. Brian 2 is another widely used network simulator that enables the definition of models directly via differential equations [4]. Furthermore, Brian 2 enables high flexibility by generating an intermediate abstract code, processed by so-called device interfaces. This allows to seamlessly exchange the underlying numerical backend architecture. While Brian 2 does not yet come with comprehensive support for multi-compartment neurons, a new extension called Dendrify specifically focuses on detailed dendritic morphology and may provide a valuable tool for the investigation of dendritic processing [85]. CARLsim [86] is another framework for the parallelized simulation of large-scale spiking neural networks, but it has been optimized for real-time simulations and neuromorphic hardware implementation. Accordingly, the single-compartment Izhikevich neuron is the most biologically realistic neuron model that is supported by CARLsim. Finally, the recently released simulator EDEN [87] appears to offer high flexibility by supporting NeuroML model descriptions, however, it still lacks support for GPU backends. As opposed to that, the equally new simulator NEST GPU (previously named NeuronGPU) [88, 89] is optimized for neural network simulations on GPU, however, it does not support multi-compartment models. Similarly, the GeNN simulator also offers highly optimized GPU simulation for networks of point neurons [90].
@@
In summary, each of the existing simulators comes with its own strengths and weaknesses (also see [91, 92, 93]). The decision to use one specific simulator should depend on the particular needs of a research project, the available compute resources, as well as the expertise of the involved researchers. For example, Brian 2 can be considered quite user-friendly for quickly setting up networks of point neurons, NEST exhibits unique performance for very large network simulations, and for NEURON there are many existing implementations of morphological neuron models. Arbor, finally, was designed to easily define networks of morphological neurons and then to map the internal modeling primitives to available compute resources in an optimized manner. As we have shown, this approach enables very good runtime efficiency and makes Arbor an attractive option, particularly, for researchers who rely on high-performance computing for their simulations with networks of morphologically detailed neurons and existing NEURON models.
@@
We have cross-validated all of our presented Arbor implementations either with Brian 2 [4] or with one of multiple stand-alone simulators that were custom-developed for previous studies [20, 59, 46]. At the single-synapse level, we did not find any significant differences between the results of the simulators (Figs. 2–5; Supplementary Figs. S1–S3). However, this does not preclude that certain differences may arise in other simulation paradigms, where the validity of the particular simulators that are used should be carefully evaluated (also cf. [94]). Furthermore, differences in the neuron model and in the numerical methods, as well as complex network effects, give rise to certain quantitative deviations in memory recall performance at the circuit level, while the qualitative behavior remained the same (Supplementary Fig. S10; also cf. subsection 3.5 and Supplementary Fig. S4).
@@
Regarding the use of compute resources for a large network with plastic connections, we found that Arbor performs well both in terms of runtime and memory use. Compared to optimized point-neuron simulators, Arbor only uses slightly more resources when computing on CPU, and can even outperform those simulators when using its capability to compute on GPU (Fig. 8). However, Arbor particularly shines when it comes to simulating networks of multi-compartment neurons, which necessitates almost no additional cost compared to single-compartment neurons (shown by the runtime per spike, Supplementary Fig. S12). This is not entirely unexpected, since Arbor has been designed particularly for networks of multi-compartment neurons. Furthermore, we found that Arbor outperforms its main competitor in this regard, CoreNEURON, in terms of runtime and memory use across all considered network sizes, even with an additional plasticity workload (Tables 8 & 9, Supplementary Table S2).
@@
The benchmarking results notwithstanding, there may be several ways to even further improve the performance of Arbor. First, to compute the dynamics of given models, Arbor uses implicit solving methods, which have the advantage of being stable but come at the cost of runtime performance. In specific cases, these methods may be replaced by faster explicit algorithms. Kobayashi et al. [95], for instance, have shown how an explicit method with adaptive time steps and second-order accuracy can serve to avoid heavy memory access, which can be helpful particularly when using GPUs. Second, the introduction of exact point neurons may eliminate the need to simulate a spatial neuron model if this is not needed. Note that meanwhile, a LIF neuron feature has been added to Arbor, although still being in a test stage.
@@
By the final model simulation results that we have presented in Fig. 6, we could gain first new insights into the functional interplay between, on the one hand, network parameters such as the size of a stimulated pattern, and on the other hand, PRP diffusion within neurons. Specifically, the results suggest that the functionality of pattern recall is not very sensitive to the diffusion speed. Nevertheless, if the diffusion is too slow, as expected, all memory functionality will vanish, which also occurs if the pattern size is too small. Moreover, if we increase the size of the dendritic branches (in a range that roughly relates to lengths of main branches in cortical pyramidal cells [77, 75, 72]), the memory recall capability is impaired (Fig. 6f–h), indicating that additional mechanisms may be needed to obtain a stable memory system. Somehow in contrast to that, a larger cell and dendrite diameter improves memory recall (Fig. 6j–l). By targeting the impact that PRP diffusion within neurons has on the dynamics of large networks, our findings may complement the picture that other studies have drawn of the functional role of spatial PRP dynamics within neurons [96, 97, 98, 99, 100].
@@
In the future, most importantly, the extended Arbor simulation framework can enable researchers to conduct studies that examine the interplay between neuron-internal, synaptic, and network dynamics. These may, for instance, be related to memory consolidation [22, 64, 101] or working memory [102, 103]. Moreover, Arbor allows the implementation of models that include changes in the connectivity structure of networks. Using this together with the tools that we have presented here will enable researchers to also study the interactions between structural and functional plasticity processes at the network level (cf. [104, 105, 106, 107, 108]). Another important goal will be to add the models that have so far been implemented in Arbor to open databases, such as the Open Source Brain repository or ModelDB, for broader adoption and community-driven validation. It would also be intriguing to conduct collaborative studies where Arbor is used to simulate electrophysiological and network properties, while at the same time using calcium imaging experiments to validate diffusion dynamics. This could serve to validate models as given in section 3.4 [45]. Furthermore, spine density distributions derived from in vivo experiments [109, 110] could be used, for example, to test diffusion models in more detail. It should further be noted that by considering synaptic plasticity processes in neural networks, simulation software contributes essentially to the understanding of biological mechanisms as well as to the development of artificial intelligence applications. In this light, further applications of our extended version of Arbor may target, for example, paradigms of somato-dendritic mismatch error reduction [111], reservoir computing with heterogeneous networks [112, 113], or prototyping and benchmarking of neuromorphic algorithms (cf. [92, 114, 115]).
@@
Due to its modern computing architecture and inherent support of multi-compartment neurons, the Arbor simulator constitutes an important tool for the computational modeling of neuronal networks. Through the newly introduced extensions, enabling to simulate synaptic plasticity, we have increased the range of Arbor’s use cases substantially. Furthermore, as an example, we have provided a first demonstration of how the extended framework enables to gain new insights into the functional impact of morphological neuronal structure at the network level. With its newly extended functionality, the Arbor framework is able to power a great variety of future studies considering synaptic mechanisms and their interactions with neuronal morphologies, from single synapses to large networks, in a highly efficient manner.
4.2 Data and code availability
@@
The Arbor framework can be installed as described on https://arbor-sim.org. The code can be retrieved, e.g., from [7]. Our simulation code for the different subsections is referenced in Table 10 below. Note that since Arbor is still under development, some parts of our model implementations may be subject to changes when using later Arbor versions, however, we intend to keep the model implementations updated along with the Arbor core code.
Due to the ongoing development of the Arbor core code, we ran our extensive simulations with different Arbor versions. For the simulation results presented in subsections 3.1–3.3 as well as 3.5–3.6, we used Arbor version 0.9.1-dev-2f4c325 (for instructions on how to install this exact version, see the README.md file that comes with our model code [68]). For the simulation results of subsection 3.4, we used the development branch of pull request #2226 with the state of commit f0e456d, which meanwhile has been merged (note that at that state of commit, Arbor still used diffusion constants in non-standard units, scaled by a factor of ). For the benchmarking results in subsection 3.7 we used Arbor version 0.10.0, and for those in subsection 3.8 we used the state of commit cebb65b of an external fork, which shall be merged upon publication of this article. Note that while all of the mentioned Arbor versions are by now slightly outdated, we ensured that our extensive software tests from [68, 116] passed with the used versions including the release version 0.10.0, which strongly indicates that the simulation results will remain the same.
@@ Acknowledgments
We would like to thank the other Arbor developers for their support and for their dedication to continuously improve the software. We would further like to thank Silvio Rizzoli, Arash Golmohammadi, and Marcel Stimberg for helpful discussions.
Financial Disclosure
For this work, CT received funding from the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) through grants SFB1286 (C01, Z01) and TE 1172/7-1, as well as from European Commission Horizon 2020 through grant no. 899265 (ADOPD). Further, CT and SSc received funding from European Commission Horizon 2020 through grant no. 945539 (HBP SGA3). The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
@@
Conceptualization: JL, SSc, CT; Data curation: JL, SSc, SSh; Formal Analysis: JL, SSc, SSh; Funding acquisition: SSc, CT; Investigation: JL, SSc, SSh, CT; Methodology: JL, SSc, SSh, TH, CT; Project administration: CT; Resources: TH, CT; Software (Arbor core): JL, SSc, TH, FB (with larger contributions by TH and FB); Software (sections 3.1–3.2): JL, SSc; Software (section 3.3): SSc, FB; Software (section 3.4): SSh; Software (sections 3.5–3.7): JL; Software (section 3.8): JL and TH; Supervision: JL, SSc, CT; Validation: JL, SSc, SSh, TH, FB; Visualization: JL, SSc, SSh, TH; Writing – original draft: JL, SSc, SSh, TH, FB; Writing – review & editing: JL, SSc, CT.